OpenStudy (anonymous):

Having a brain fart with domain and range again.. Just getting throw off because the x squared and x. Please little help. Dont have to give the answer, just how to get it. Its. Square root of x^2-3x

OpenStudy (anonymous):

\[x^2-3x \ge0\]

OpenStudy (anonymous):

take the argument of the function, what is inside the radical and set in an inequality where the argument has to be greater of equal to zero. \[x^2-3x \ge0\]

OpenStudy (anonymous):

solver for x and then you will get what x cannot equal to. so your domain would be all reals except what ever x is equal to.

OpenStudy (anonymous):

I got that far giggles but now not sure because the 2 xs

OpenStudy (anonymous):

domain is real number R because any polynomila its domain would be real number inless it's mentioned

OpenStudy (anonymous):

Would it be x(x-3)>0. so X can not be less then or equal to 3?

OpenStudy (anonymous):

But no because can be negative..

OpenStudy (anonymous):

factor, find the zeros, graph it will be obvious where it is greater than 0

OpenStudy (anonymous):

I have to solve with out graphing. Range and domain

OpenStudy (anonymous):

{x ge\[x \ge 3\]

OpenStudy (anonymous):



OpenStudy (anonymous):

for the range , this funcction represent a parabola with a vertex (-b/a, f(-b/a)) = (3/2,-9/4) ===> the range =[-9/4, +infinity)

OpenStudy (anonymous):

i didn't mean graph the function itself, i meant graph \[y=x^2-3x=x(x-3)\] get a parabola that opens up positive outside the zeros, negative between them

OpenStudy (anonymous):

Wont the domain be X>3 and X<0?

OpenStudy (anonymous):

the function is \[f(x)=\sqrt{x^2-3x}\] right? so you have to make sure \[x^2-3x\geq 0\]

OpenStudy (anonymous):

yes that is the domain you have it

OpenStudy (anonymous):

Yes satellite. But how do i get that dominate without graphing? My equation only gets >3

OpenStudy (anonymous):

\[x\leq 0\text{ or } x\geq 3\]

OpenStudy (anonymous):

you mean the inequality \[x(x-3)\geq 0\]?

OpenStudy (anonymous):

domain not dominate sorry

OpenStudy (anonymous):

without graphing you say "if x > 3 both factors are positive, so their product is positive and if x < 0 both factors are negative so their product is positive

OpenStudy (anonymous):

I get that inequality. But when i solve it how am I finding the x<0

OpenStudy (anonymous):

what i wrote, although the parabola in your head should get it as well. you can just reason it out. there is no need for algebra

OpenStudy (anonymous):

if \[x<0\] than both x and (x-3) are negative so their product is positive

OpenStudy (anonymous):

The problem is my professor wants the algebra.. I know how to get it graphing obviously

OpenStudy (anonymous):

well it really isn't algebra is it? you can make a little sign chart if you like

OpenStudy (anonymous):

x ------------- 0 ++++++++++++++++++ x-3 -------------------------- 3 ++++++++++ x(x-3) +++++++0 ---------- 3 ++++++++++

OpenStudy (anonymous):

As I stated above domain is clearly the whole real number set R because our function is a polynomial. and the range is (-9/4,+infinity)

OpenStudy (anonymous):

kind of childish though don't you think? you already know that a parabola that faces up is positive outside the zeros and negative between them

OpenStudy (anonymous):

@raheen it is \[f(x)=\sqrt{x^2-3x}\] not \[f(x)=x^2-3x\]

OpenStudy (anonymous):

btw range is \[[0,\infty)\] because square root symbol means the positive root

OpenStudy (anonymous):

oh . I missed the square root notation, I'm sorry.

OpenStudy (anonymous):

you are talking about what been posted but I'm solving for different data satellite

OpenStudy (anonymous):

I dont think think thats the range. When i graph it, doesnt hit 0

OpenStudy (anonymous):

domain is Real Number except (0,3)

OpenStudy (anonymous):

sure it does. \[f(0)=0 \] right?

OpenStudy (anonymous):

You mean 0 to 3?

OpenStudy (anonymous):

in fact \[f(3)=0\] as well

OpenStudy (anonymous):

domain \[(-\infty,0]\cup [3,\infty)\] range \[[0,\infty)\]

OpenStudy (anonymous):

Alright. Just wish I knew how to solve without graphing

OpenStudy (anonymous):

hold on. what are you trying to solve "without graphing" the range?